OpenStudy (anonymous):

Write an equation of the line that passes through the point (-4,-1) and is (a) parallel to and (b) perpendicular to the line 2x + 7y = 14.

OpenStudy (bibby):

re arrange the line 2x + 7y = 14. 7y = -2x+14 y = (-2/7)x+2 you know your slop is going to be the negative reciprocal of the slope in that line

OpenStudy (anonymous):

Is the first answer to parallel

OpenStudy (bibby):

the first answer has the slope -2/7 but the equation looks different

OpenStudy (anonymous):

Okay but is the first answer the answer to (a)?

OpenStudy (bibby):

y=mx+b -1=7/2 * -4+b b=-1-(7/2)(-4). b=13. y=7/2x+13

OpenStudy (anonymous):

Alright, is that the parallel?

OpenStudy (bibby):

no. look at the slopes. that is the negative reciprocal of the slope given. Perpendicular

OpenStudy (bibby):

\[y= \frac{ -2 }{ 7? }x-\frac{ 15 }{ 7 }\]

OpenStudy (anonymous):

Okay well this means....I'm one step closer to finishing my time machine. Thank you random citizen.

OpenStudy (bibby):

I love you too